# Vision Transformers in Precision Agriculture: A Comprehensive Survey

### v2

@@ 4.2 Hybrid Models
In a comparative study, the authors evaluate the performance of standalone and hybrid models and propose a lightweight DL approach utilizing ViT for automated plant disease classification [9]. They compare ViT-based models, CNNs, and hybrid architectures that combine both techniques. These models were evaluated on three different datasets: the Wheat Rust Classification Dataset (WRCD), the Rice Leaf Disease Dataset (RLDD), and the PlantVillage dataset. A ViT-based model achieved a 100% F1-score on WRCD with a 200×200 resolution, outperforming EfficientNetB0 in computational efficiency. For RLDD, hybrid models balanced accuracy and computational cost, with the ViT-based model achieving the highest average precision. On the larger PlantVillage dataset, the ViT-only model reached an F1-score of 98.77%, surpassing CNN-based architectures. Models that combine attention blocks and CNNs as hybrid models offered a balance between accuracy and speed. They were comparatively faster than pure ViT models and more accurate than CNN-only models. For instance, a hybrid model achieved high accuracy in the PlantVillage experiments, indicating the potential of integrating attention mechanisms into CNNs for real-world applications. It is worth mentioning that this study highlights the need for a practical, lightweight, real-time implementation for agricultural purposes. Multi-label classification with object localization for detailed disease identification is the focus of future research.
EfficientRMT-Net is a novel hybrid model that integrated ResNet-50 and ViT in order to classify potato plant leaf diseases effectively [70]. EfficientRMT-Net employs depth-wise convolution for computational efficiency and utilizes a stage-block architecture to enhance scalability and feature sensitivity. This paper emphasizes that the ability of EfficientRMT-Net to process high-dimensional data efficiently while maintaining high accuracy positions it as a valuable tool for smart farming. [95] proposed SEViT model combining Squeeze-and-Excitation ResNet101 and ViT to tackle large-scale and fine-grained plant disease classification challenges. The preprocessing network employs Squeeze-and-Excitation modules in ResNet101 to enhance channel-specific feature representation, while the ViT component leverages its global attention mechanism for feature extraction and classification. SEViT improves upon the limitations of CNN-based methods, which often struggle to distinguish diseases in visually similar crops. Although comparisons with popular models such as MobileNetV3, EfficientNet, and VGG16 highlight the efficiency of SEViT in handling visually similar diseases across different crops, SEViT faces challenges including high computational requirements due to the depth of the preprocessing network and its reliance on pretrained weights to optimize the performance of ViT.
@@
A novel approach using ViTs was also introduced to address challenges in paddy leaf disease identification [68]. The model incorporates a multi-scale contextual feature extraction mechanism to capture both local and global representations of lesions on paddy leaves. This is complemented by a weakly supervised Paddy Lesion Localization (PLL) unit that prioritizes significant lesion areas to enhance classification accuracy. The model’s performance was evaluated on the Paddy Doctor dataset [58]. The proposed model outperformed several state-of-the-art models, including DenseNet169, ResNet50, and SwinT, achieving an impressive accuracy of 98.74%, an F1-score of 98.18%, and an AUC of 99.49%. These results demonstrate significant improvements over baseline models, highlighting the efficacy of lesion-focused feature extraction and context-aware learning in accurately classifying complex diseases. The authors emphasize that integrating lesion localization with multi-scale ViT structures enables robust performance across diverse visual challenges. The model’s enhanced ability to identify and prioritize discriminative lesion areas provides more reliable disease detection, supporting its potential for real-world applications in precision agriculture.
@@
In another study, the authors introduced an ensemble learning approach with hard and soft voting strategies that integrates three CNN models (MobileNetV3, DenseNet201, ResNext50) and two Transformer models (ViT and SwinT) to classify leaf diseases [42]. Experimental results show that ensemble learning improved classification accuracy, with ViT achieving superior performance as a standalone model. The authors also employed Grad-CAM visualization to highlight the regions of input images most influential for classification decisions, demonstrating that the models effectively localized disease areas. Their results show that ensemble learning combining CNN and Transformer models provides a powerful approach for effective classification of plant diseases, which is useful in practical agriculture.
@@
A novel ensemble model, Residual Swin Transformer Networks (RST-Nets), was proposed to address the challenges of plant disease recognition, including noise, varying image intensities, and the subtle differences between healthy and diseased plants [36]. The model integrates residual convolutional blocks with SwinT, leveraging the latter’s hierarchical architecture for scalable complexity and global context awareness. Residual connections enhance the feature extraction process by retaining crucial information from earlier layers. The model achieved outstanding results, surpassing state-of-the-art models such as ResNet and GoogleNet, but its computational overhead limited its suitability for resource-constrained IoT and edge devices.
@@
Another way to create a hybrid model is to build a multi-stage model that combines CNN and ViT by dividing the network into sequential stages. This hierarchical approach reduces computational costs while preserving essential information, enabling robust performance in complex environments. As shown in Figure 8, [45] proposed a novel model, ConvViT, which combines ViT and CNNs to identify kiwifruit diseases in complex natural environments. The authors developed a custom dataset of 25,168 images capturing six types of kiwifruit leaf diseases and validated the model’s performance on this dataset, as well as on the publicly available PlantVillage and AIChallenger2018 datasets. To enhance feature extraction, the model employs overlapping patch embeddings and alternating convolutional and Transformer layers, ensuring both global and local feature representation. On the kiwifruit dataset, ConvViT achieved a top identification accuracy of 98.78%, surpassing benchmarks like ResNet, ViT, and ResMLP, while maintaining a lightweight design with reduced parameters and FLOPs. Key improvements include overlapping patch embedding to preserve local image continuity and reduce computational complexity, as well as a multi-stage design to balance feature extraction efficiency and computational cost. The Transformer’s global attention complements CNN’s local feature extraction, making ConvViT highly effective in complex environments with variable lighting, backgrounds, and noise. The model’s innovative design and lightweight adaptability make it a valuable backbone for broader identification tasks in agriculture, potentially aiding real-world crop disease management efforts.
@@
There are also studies that use transfer learning in addition to hybrid models. In the research by [78], the authors propose a hybrid model, TLMViT, for plant disease classification that combines transfer learning models followed by a ViT. They incorporate data augmentation techniques, which address class imbalance and mitigate overfitting, along with a two-phase feature extraction process: initial extraction using pre-trained CNNs models (e.g., VGG16, VGG19, ResNet50, AlexNet, and Inception V3), and deep feature extraction using ViT. Final classification is conducted using an MLP classifier. The model achieved a validation accuracy of 98.81% on the PlantVillage dataset (VGG19-ViT combination) and 99.86% on the Wheat Rust dataset. These results represent a 1.11% and 1.099% improvement in accuracy compared to transfer learning models without ViT. These findings indicate the effectiveness of combining transfer learning and ViT for deep feature extraction. The authors also highlight that TLMViT effectively leverages pre-trained models to reduce dimensionality and computational complexity.
@@
There are some studies that use multi-track—especially dual-track (or dual-path)—models that integrate two parallel architectures, for example, a ViT and a CNN, to leverage the strengths of both. Figure 9 shows COFFORMER, a novel dual-path visual Transformer designed for efficient and interpretable diagnosis of coffee leaf diseases, which integrates lesion segmentation and disease classification into a unified framework [67]. The segmentation path employs a U-shaped architecture built on COFFORMER blocks, which use multiscale convolutional pooling for token mixing, while the classification path incorporates a Coffee Lesion Attention (CLA) module to focus on critical lesion regions. The framework evaluates both the type and severity of diseases, enabling real-time and explainable disease diagnosis. The dataset used for evaluation consists of 1,685 images of Arabica coffee leaves captured under controlled conditions and annotated for five disease types and severity levels. Experimental results demonstrate COFFORMER’s superior performance compared to state-of-the-art models. For segmentation, it achieved a Dice Similarity Coefficient (DSC) of 98.14% and a mean Intersection over Union (mIoU) of 97.98%, outperforming baselines such as Swin-Unet and PSPNet. In classification tasks, COFFORMER achieved an accuracy of 99.1% and an F1-score of 99.32% for disease type detection, alongside 99.12% accuracy for severity estimation. These results demonstrate its robustness in identifying diverse lesions and disease severity across real-world coffee leaf images.
@@
A real-time plant disease identification framework called CondConViT has been proposed, integrating a Vision Transformer (ViT) with a Conditional Convolutional Neural Network (CondConv), and further enhanced by a novel attention mechanism—Conditional Attention with Statistical Squeeze-and-Excitation (CASSE) [80]. To improve generalization and robustness, the authors introduce a data augmentation technique based on a modified CycleGAN, termed C3GAN, which synthesizes realistic in-field images from healthy plant samples. The model is designed to be lightweight, with only 0.95 million parameters, and is deployable on edge devices for drone-based surveillance. CondConViT is evaluated against seven state-of-the-art models and demonstrates superior interpretability and classification performance across diverse environmental conditions using both Grad-CAM and LIME techniques. They used six datasets: five public datasets—PlantVillage [51], Embrapa [5], Plantpathology [82], Maize and Rice [14], and a newly created in-field drone image dataset named IIITDMJ-Maize [81], which includes 416 labeled images of maize plants under various weather conditions. C3GAN was employed to augment the IIITDMJ-Maize dataset by generating 300 synthetic images for three maize diseases: common rust, northern leaf blight, and gray leaf spot. The CondConViT model achieved the highest or near-highest performance across all datasets in terms of accuracy, F1-score, and AUC. Furthermore, it showed strong generalization on unseen raw drone images and outperformed heavier models, making it suitable for real-time and resource-constrained agricultural applications.
@@ 5.1 Key Findings
Several important findings have emerged from the review of ViTs in precision agriculture. ViTs have shown substantial improvements over traditional CNNs in various agricultural tasks, notably in plant disease detection and crop monitoring. Their ability to capture long-range relationships within images enables more accurate identification of plant diseases and pests, even in complex environments. Combining ViTs with CNNs in hybrid models has further boosted classification accuracy and robustness, making them a promising approach for agricultural challenges. Furthermore, transfer learning—where ViTs pre-trained on large datasets are fine-tuned for agricultural applications—has proven effective in addressing data limitations, improving model performance without requiring vast amounts of labeled data. We can summarize the key findings as follows:
@@ 5.2 Open Challenges
  • Need for more comprehensive and diverse datasets: Many studies use publicly available datasets such as PlantVillage, which may not fully represent the variability of real-world scenarios. Real-world conditions involve a multitude of factors, such as diverse plant species, various disease types and stages, and different environmental and lighting conditions—factors that are often not captured in standard datasets. Datasets also need to account for image variations, including different angles, distances, and image quality. The lack of diverse data can lead to models that are not robust and may not generalize well to new, unseen data. Although some studies have created their own datasets to address this issue, collecting and annotating large-scale, high-quality datasets remains expensive and time-consuming. Table LABEL:tab:tabledata shows the number of times each dataset has been used.
@@
  • Reducing computational costs: Transformer-based models can be computationally expensive, requiring significant computing resources. This cost stems from the self-attention mechanism, whose complexity scales with the input size, making it more difficult to train on larger images and datasets. As a result, deploying these models on resource-constrained devices—such as smartphones or edge devices—can be challenging, especially in areas with limited access to high-performance computing infrastructure. Some studies propose solutions such as pruning less important attention heads or using sparse matrix multiplication to reduce computational overhead and enhance practicality in real-world applications. Other approaches, including knowledge distillation and hybrid models that combine ViTs with CNNs, also aim to improve the efficiency of ViTs for deployment. Further research is needed to develop more efficient architectures and training techniques to make ViTs more accessible and suitable for practical use in plant disease detection.
@@
PlantVillage [51] Contains over 50,000 images of plant leaves across 14 species and 26 diseases $18$
@@
From Kaggle Other publicly available datasets $12$
Custom Datasets gathered from search engines, created by combining other datasets, or not easily accessible $8$
PlantPathology2020 [82] Contains over 3,000 high-quality real-life images of apple leaf diseases under varying conditions $3$
PlantDoc [76] Contains over 2,000 images across 13 plant species and 17 disease classes $2$
@@
Embrapa [5] Contains over 56,000 images of 171 diseases and other disorders affecting 21 plant species $1$

### v3

@@ 4.2 Hybrid Models
In a comparative study, the authors evaluate the performance of standalone and hybrid models and propose a lightweight DL approach utilizing ViT for automated plant disease classification [9]. They compare ViT-based models, CNNs, and hybrid architectures that combine both techniques. These models were evaluated on three different datasets: the Wheat Rust Classification Dataset (WRCD), the Rice Leaf Disease Dataset (RLDD), and the PlantVillage dataset. A ViT-based model achieved a 100% F1-score on WRCD with a 200×200 resolution, outperforming EfficientNetB0 [93] in computational efficiency. For RLDD, hybrid models balanced accuracy and computational cost, with the ViT-based model achieving the highest average precision. On the larger PlantVillage dataset, the ViT-only model reached an F1-score of 98.77%, surpassing CNN-based architectures. Models that combine attention blocks and CNNs as hybrid models offered a balance between accuracy and speed. They were comparatively faster than pure ViT models and more accurate than CNN-only models. For instance, a hybrid model achieved high accuracy in the PlantVillage experiments, indicating the potential of integrating attention mechanisms into CNNs for real-world applications. It is worth mentioning that this study highlights the need for a practical, lightweight, real-time implementation for agricultural purposes. Multi-label classification with object localization for detailed disease identification is the focus of future research.
EfficientRMT-Net [82] is a novel hybrid model that integrated ResNet-50 and ViT in order to classify potato plant leaf diseases effectively. EfficientRMT-Net employs depth-wise convolution for computational efficiency and utilizes a stage-block architecture to enhance scalability and feature sensitivity. This paper emphasizes that the ability of EfficientRMT-Net to process high-dimensional data efficiently while maintaining high accuracy positions it as a valuable tool for smart farming. [114] proposed SEViT model combining Squeeze-and-Excitation [40] ResNet101 and ViT to tackle large-scale and fine-grained plant disease classification challenges. The preprocessing network employs Squeeze-and-Excitation modules in ResNet101 to enhance channel-specific feature representation, while the ViT component leverages its global attention mechanism for feature extraction and classification. SEViT improves upon the limitations of CNN-based methods, which often struggle to distinguish diseases in visually similar crops. Although comparisons with popular models such as MobileNetV3 [37], EfficientNet, and VGG16 highlight the efficiency of SEViT in handling visually similar diseases across different crops, SEViT faces challenges including high computational requirements due to the depth of the preprocessing network and its reliance on pretrained weights to optimize the performance of ViT.
@@
A novel approach using ViTs was also introduced to address challenges in paddy leaf disease identification [78]. The model incorporates a multi-scale contextual feature extraction mechanism to capture both local and global representations of lesions on paddy leaves. This is complemented by a weakly supervised Paddy Lesion Localization (PLL) unit that prioritizes significant lesion areas to enhance classification accuracy. The model’s performance was evaluated on the Paddy Doctor dataset [67]. The proposed model outperformed several state-of-the-art models, including DenseNet169 [41], ResNet50, and SwinT, achieving an impressive accuracy of 98.74%, an F1-score of 98.18%, and an AUC of 99.49%. These results demonstrate significant improvements over baseline models, highlighting the efficacy of lesion-focused feature extraction and context-aware learning in accurately classifying complex diseases. The authors emphasize that integrating lesion localization with multi-scale ViT structures enables robust performance across diverse visual challenges. The model’s enhanced ability to identify and prioritize discriminative lesion areas provides more reliable disease detection, supporting its potential for real-world applications in precision agriculture.
@@
In another study, the authors introduced an ensemble learning approach with hard and soft voting strategies that integrates three CNN models (MobileNetV3, DenseNet201, ResNext50 [109]) and two Transformer models (ViT and SwinT) to classify leaf diseases [51]. Experimental results show that ensemble learning improved classification accuracy, with ViT achieving superior performance as a standalone model. The authors also employed Grad-CAM visualization to highlight the regions of input images most influential for classification decisions, demonstrating that the models effectively localized disease areas. Their results show that ensemble learning combining CNN and Transformer models provides a powerful approach for effective classification of plant diseases, which is useful in practical agriculture.
@@
A novel ensemble model, Residual Swin Transformer Networks (RST-Nets), was proposed to address the challenges of plant disease recognition, including noise, varying image intensities, and the subtle differences between healthy and diseased plants [44]. The model integrates residual convolutional blocks with SwinT, leveraging the latter’s hierarchical architecture for scalable complexity and global context awareness. Residual connections enhance the feature extraction process by retaining crucial information from earlier layers. The model achieved outstanding results, surpassing state-of-the-art models such as ResNet and GoogleNet [91], but its computational overhead limited its suitability for resource-constrained IoT and edge devices.
@@
Another way to create a hybrid model is to build a multi-stage model that combines CNN and ViT by dividing the network into sequential stages. This hierarchical approach reduces computational costs while preserving essential information, enabling robust performance in complex environments. As shown in Figure 8, [54] proposed a novel model, ConvViT, which combines ViT and CNNs to identify kiwifruit diseases in complex natural environments. The authors developed a custom dataset of 25,168 images capturing six types of kiwifruit leaf diseases and validated the model’s performance on this dataset, as well as on the publicly available PlantVillage and AIChallenger2018 datasets. To enhance feature extraction, the model employs overlapping patch embeddings and alternating convolutional and Transformer layers, ensuring both global and local feature representation. On the kiwifruit dataset, ConvViT achieved a top identification accuracy of 98.78%, surpassing benchmarks like ResNet, ViT, and ResMLP [100], while maintaining a lightweight design with reduced parameters and FLOPs. Key improvements include overlapping patch embedding to preserve local image continuity and reduce computational complexity, as well as a multi-stage design to balance feature extraction efficiency and computational cost. The Transformer’s global attention complements CNN’s local feature extraction, making ConvViT highly effective in complex environments with variable lighting, backgrounds, and noise. The model’s innovative design and lightweight adaptability make it a valuable backbone for broader identification tasks in agriculture, potentially aiding real-world crop disease management efforts.
@@
There are also studies that use transfer learning in addition to hybrid models. In the research by [92], the authors propose a hybrid model, TLMViT, for plant disease classification that combines transfer learning models followed by a ViT. They incorporate data augmentation techniques, which address class imbalance and mitigate overfitting, along with a two-phase feature extraction process: initial extraction using pre-trained CNNs models (e.g., VGG16, VGG19, ResNet50, AlexNet [48], and Inception V3), and deep feature extraction using ViT. Final classification is conducted using an MLP classifier. The model achieved a validation accuracy of 98.81% on the PlantVillage dataset (VGG19-ViT combination) and 99.86% on the Wheat Rust dataset. These results represent a 1.11% and 1.099% improvement in accuracy compared to transfer learning models without ViT. These findings indicate the effectiveness of combining transfer learning and ViT for deep feature extraction. The authors also highlight that TLMViT effectively leverages pre-trained models to reduce dimensionality and computational complexity.
@@
Some studies utilize multi-track models, particularly dual-track (or dual-path) architectures, that integrate two parallel networks, such as a ViT and a CNN, to leverage the strengths of both. Figure 9 shows COFFORMER, a novel dual-path visual Transformer designed for efficient and interpretable diagnosis of coffee leaf diseases, which integrates lesion segmentation and disease classification into a unified framework [77]. The segmentation path employs a U-shaped architecture built on COFFORMER blocks, which use multiscale convolutional pooling for token mixing, while the classification path incorporates a Coffee Lesion Attention (CLA) module to focus on critical lesion regions. The framework evaluates both the type and severity of diseases, enabling real-time and explainable disease diagnosis. The dataset used for evaluation consists of 1,685 images of Arabica coffee leaves captured under controlled conditions and annotated for five disease types and severity levels. Experimental results demonstrate COFFORMER’s superior performance compared to state-of-the-art models. For segmentation, it achieved a Dice Similarity Coefficient (DSC) of 98.14% and a mean Intersection over Union (mIoU) of 97.98%, outperforming baselines such as Swin-Unet [12] and PSPNet [118]. In classification tasks, COFFORMER achieved an accuracy of 99.1% and an F1-score of 99.32% for disease type detection, alongside 99.12% accuracy for severity estimation. These results demonstrate its robustness in identifying diverse lesions and disease severity across real-world coffee leaf images.
@@
A real-time plant disease identification framework called CondConViT has been proposed, which integrates a ViT with a Conditional Convolutional Neural Network (CondConv) and is further enhanced by a novel attention mechanism known as Conditional Attention with Statistical Squeeze-and-Excitation (CASSE). [96]. To improve generalization and robustness, the authors introduce a data augmentation technique based on a modified CycleGAN, termed C3GAN, which synthesizes realistic in-field images from healthy plant samples. The model is designed to be lightweight, with only 0.95 million parameters, and is deployable on edge devices for drone-based surveillance. CondConViT is evaluated against seven state-of-the-art models and demonstrates superior interpretability and classification performance across diverse environmental conditions using both Grad-CAM and LIME techniques. They used six datasets, including five public datasets: PlantVillage [60], Embrapa [5], Plantpathology [99], Maize and Rice [15], and a newly created in-field drone image dataset named IIITDMJ-Maize [97], which comprises 416 labeled images of maize plants captured under various weather conditions. C3GAN was employed to augment the IIITDMJ-Maize dataset by generating 300 synthetic images for three maize diseases: common rust, northern leaf blight, and gray leaf spot. The CondConViT model achieved the highest or near-highest performance across all datasets in terms of accuracy, F1-score, and AUC. Furthermore, it showed strong generalization on unseen raw drone images and outperformed heavier models, making it suitable for real-time and resource-constrained agricultural applications.
In another study [98], a parallel fusion of CNN and ViT modules was proposed to capture both local and global features. Unlike sequential CNN-ViT hybrids, this approach enables independent feature extraction followed by fusion using MobileNetV2 [79] and Squeeze-and-Excitation (SE) blocks, thereby improving efficiency and interpretability. ConViTX is ultra-lightweight, comprising only 704,882 trainable parameters and requiring 0.647 GFLOPs, making it ideal for deployment on resource-constrained IoT/edge devices. Explainability is achieved using Grad-CAM and LIME, and the model’s feature discrimination capability is further validated through t-SNE visualizations. The model was rigorously evaluated on five datasets: PlantVillage, Embrapa, PlantDoc, a custom in-field maize dataset (IIITDM-Maize), and a combined multi-species dataset called PlantCOMB. On PlantVillage, ConViTX achieved an accuracy of 99.63%, an F1-score of 99.64%, and an AUC of 99.96%, outperforming nine state-of-the-art (SOTA) models. On Embrapa, it achieved an F1-score of 95.12% and an AUC of 99.74%, and on PlantDoc, it retained the top spot in most metrics. On the custom maize dataset, it achieved 93.02% accuracy, matching the best competitors while showing a 2% higher AUC. On raw drone images, it reached 61.42% accuracy, surpassing most SOTA models in generalizability. Beyond static evaluation, ConViTX was deployed on a Jetson Nano-equipped drone and tested on real-time video frames captured from maize farms. It processed frames at approximately 29 FPS and demonstrated competitive results, notably identifying northern leaf blight with 58% accuracy and healthy plants with 73% accuracy. Though challenges remain with subtle diseases like common rust, the model achieved a better balance between speed, accuracy, and interpretability compared to other lightweight and heavyweight models. Overall, ConViTX proves to be an efficient, interpretable, and deployable solution, opening a promising avenue for hybrid CNN-ViT architectures in real-world smart agriculture systems.
RepAgrViT [18], a novel lightweight hybrid architecture, is proposed to integrate CNN and ViT principles for plant disease recognition in Internet of Things (IoT) edge environments. Unlike traditional CNN-ViT hybrids, RepAgrViT employs a unique dual-stream design consisting of a token mixer and a channel mixer arranged in series to simultaneously process local and global features. A core innovation is the Bilinear Attention Transformation (BATransform) module, which captures long-range dependencies between diseased and healthy leaf regions. Additionally, the model integrates structural reparameterization techniques to reduce inference complexity while preserving accuracy. At just 9.6M parameters, the model demonstrates strong computational efficiency while retaining competitive performance.
To optimize hyperparameters efficiently, the authors introduce Lite-AVPSO, a novel Particle Swarm Optimization variant that incorporates adaptive weighted delayed velocity and neighborhood-based local search mechanisms. This approach significantly enhances the hyperparameter search process over alternatives like TPE, GA, and Hyperband. Through an extensive optimization process spanning over 200 trials, the model achieved 98.54% average accuracy across three plant disease datasets: the Coffee Leaf Dataset , the Plant Pathology Dataset , and the Rice Leaf Disease Dataset. On these datasets, RepAgrViT achieved 98.03%, 96.99%, and 98.78% accuracy respectively. Overall, this work contributes a highly efficient and interpretable AI system tailored for precision agriculture in resource-constrained settings.
@@
[98] 5 Public Varies by dataset Varies by dataset Varies by dataset Varies by dataset Proposing an ultra-lightweight and interpretable hybrid model that combines CNN and ViT in a parallel architecture for accurate and efficient plant disease classification, suitable for deployment on resource-constrained IoT devices
[18] 5 Public , 1 Custom Varies by dataset Varies by dataset Varies by dataset Varies by dataset Proposing a novel lightweight hybrid architecture that combines CNN and ViT components through a dual-stream token and channel mixer, along with a Bilinear Attention Transformation module for efficient local-global feature extraction
@@ 5.1 Key Findings
Several important findings have emerged from the review of ViTs in precision agriculture. ViTs have shown substantial improvements over traditional CNNs in various agricultural tasks, notably in plant disease detection and crop monitoring. Their ability to capture long-range relationships within images enables more accurate identification of plant diseases and pests, even in complex environments. Combining ViTs with CNNs in hybrid models has further boosted classification accuracy and robustness, making them a promising approach for agricultural challenges. Furthermore, transfer learning, in which ViTs pre-trained on large datasets are fine-tuned for agricultural applications, has proven effective in addressing data limitations and improving model performance without requiring vast amounts of labeled data. We can summarize the key findings as follows:
@@ 5.2 Open Challenges
  • Need for more comprehensive and diverse datasets: Many studies use publicly available datasets such as PlantVillage, which may not fully represent the variability of real-world scenarios. Real-world conditions involve a multitude of factors, including diverse plant species, various disease types and stages, and different environmental and lighting conditions, which are often not captured in standard datasets. Datasets also need to account for image variations, including different angles, distances, and image quality. The lack of diverse data can lead to models that are not robust and may not generalize well to new, unseen data. Although some studies have created their own datasets to address this issue, collecting and annotating large-scale, high-quality datasets remains expensive and time-consuming. Table LABEL:tab:tabledata shows the number of times each dataset has been used.
@@
  • Reducing computational costs: Transformer-based models can be computationally expensive, requiring significant computing resources. This cost stems from the self-attention mechanism, whose complexity scales with the input size, making it more difficult to train on larger images and datasets. As a result, deploying these models on resource-constrained devices, such as smartphones or edge devices, can be challenging, particularly in areas with limited access to high-performance computing infrastructure. Some studies propose solutions such as pruning less important attention heads or using sparse matrix multiplication to reduce computational overhead and enhance practicality in real-world applications. Other approaches, including knowledge distillation and hybrid models that combine ViTs with CNNs, also aim to improve the efficiency of ViTs for deployment. Further research is needed to develop more efficient architectures and training techniques to make ViTs more accessible and suitable for practical use in plant disease detection.
@@
PlantVillage [60] Contains over 50,000 images of plant leaves across 14 species and 26 diseases $20$
From Kaggle Other publicly available datasets $14$
@@
Custom Datasets gathered from search engines, created by combining other datasets, or not easily accessible $10$
PlantPathology2020 [99] Contains over 3,000 high-quality real-life images of apple leaf diseases under varying conditions $4$
PlantDoc [89] Contains over 2,000 images across 13 plant species and 17 disease classes $4$
@@
Embrapa [5] Contains over 56,000 images of 171 diseases and other disorders affecting 21 plant species $2$